[Brahms-l] Shift report 20050607 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 07 2005 - 07:59:00 EDT
Shifter leader was sl
Shifters were gl


1. Long store to begin with. Second store lost after approximately 2 hours. Latest store lost before any data were possible.

2. 5 hrs worth of runs at 90 deg B full, and 3.0 deg full A.

3. No new problems noticed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 7 07:59:11 2005

This archive was generated by hypermail 2.1.8 : Tue Jun 07 2005 - 07:59:23 EDT